OpenStudy (lilai3):

@UnkleRhaukus Nabuko practices the violin at least 12 hours per week. She practices for three-fourths of an hour each session. If Nabuko has already practiced 3 hours in one week, how many sessions remain to meet or exceed her weekly practice goal? Please help. Explain. Thank you.~

OpenStudy (unklerhaukus):

set x as the number of sessions remaining

OpenStudy (lilai3):

okay. then...?

OpenStudy (unklerhaukus):

[goal] ≤ [hours already practiced]+[hours of practice remaining] [goal] ≤ [hours already practiced]+[hours per session]×[number sessions remaining]

OpenStudy (lilai3):

wait.. what? this is confusing..

OpenStudy (lilai3):

her goal is 12 hours?

OpenStudy (unklerhaukus):

yes

OpenStudy (lilai3):

\[12\le 3+(12-3)\]

OpenStudy (unklerhaukus):

[option] [,] \(\to\) ≤

OpenStudy (lilai3):

wait what? where's option?

OpenStudy (unklerhaukus):

[option/alt]

OpenStudy (lilai3):

no. where is option? is it on the keyboard, or right-click on the mouse or what? you are making me confused...

OpenStudy (unklerhaukus):

my keyboard has a key with 'alt' and 'option' on it

OpenStudy (lilai3):

oh, i see....okay, any way is my equation right? please scroll up to see my equation.

OpenStudy (unklerhaukus):

\(12≤3+\)[hours per session]\(×\)[number sessions remaining]

OpenStudy (lilai3):

\[12\le3+12\times(12-3)\] right?

OpenStudy (unklerhaukus):

nope,

OpenStudy (lilai3):

i don't get it then. because that is what you told me to do.

OpenStudy (unklerhaukus):

'She practices for three-fourths of an hour each session. ' [hours per session] = 3/4 [number sessions remaining] = x

OpenStudy (lilai3):

oh...isn't that just 45 minutes?

OpenStudy (unklerhaukus):

the question is asking for the number of session that remain

OpenStudy (lilai3):

okay.

OpenStudy (lilai3):

....???

OpenStudy (lilai3):

so what is the answer, because it is really late at night here, so can you please help me hurriedly? thanks at least please give me the equation? that would help GREATLY also medal be rewarded, thank you.

OpenStudy (unklerhaukus):

\[12≤3+(3/4)x\]

OpenStudy (lilai3):

thank you. is it n is greater than or equal to 12? is that the answer? because that is the answer that i got. thank you. i understand you much better now, than ever before. thank you. you are a great help.

OpenStudy (lilai3):



OpenStudy (lilai3):

thank you very much. it may not be 12:00 am in the morning at night at your place, but it is in my place, so good night. thank you EVER so much!

OpenStudy (unklerhaukus):

yeah 12 twelve sessions, because 3[hour]+(3/4)[session/hour]*12[sessions]= 12[hr]

OpenStudy (lilai3):

okay thank you. bye (please don't reply because i got to go now. bye!) bye!